3. Select a counterexample to show that the statement is false. When an odd number is added to the product of two even numbers, the result will be even.
Answer by jim_thompson5910(35256)

When an odd number is added to the product of two even numbers

Example:

3 + 2 * 4 = 3+8 = 11

which is NOT even. So the statement is false.

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------